Session: Feb 2017 (session 1)
Wavelength: 0.7 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Angel Vicente at Mon Mar 13 04:36:58 2017) Minor failures:
Not record first scan

Effelsberg (Uwe at Mon Mar 13 09:51:04 2017) Success:
No problems, beside the broken LCP channel.

Onsala (Jun Yang at Wed Mar 15 13:26:38 2017) Success:
No known problems.